16 April 2012 LUCI/LBC checkout

The science prep night last night was lost to weather. Therefore, the tasks to be done have been carried over to tonight:
  • collect data to compute the AGW3-LUCI transformation matrix [Doug, Jesper, Dave]
    • Dave left LUCI at N3.75 camera and with BrGamma filter.
    • Doug's scripts will take data
  • test transformation by offsetting in a dice5 and/or along the slit.
  • LUCI database needs to be restored [Kai]
  • LUCI camera wheel problem must be addressed.
  • LBC problems with comm dropouts. HK Blue side dropping out. Also filters/rotator comm dropouts.
    • this afternoon, I experienced the same problems. At 17:53 MST, I rebooted CMU. After this so far, so good.
      • 10 biases completed OK
      • Calib_OBs/FILTERtesting/testfiltersBR_dome

  • LBC testing:
    • testfiltersBR_dome running but LBC restarted after crash message at 01:15 UT. I see no reason in the log, no error or warning, that explains why this occurred.
    • back up and running at 01:19 and will resend OB.
    • completed OK. Try again..blue filters hardware error, but now running.
378  2012/04/16   01:47:07.231093         hinibit counter updated to 2
377   2012/04/16   01:47:07.231012         hinibit counter updated to 1
376   2012/04/16   01:47:00.910545   FILTERS    WHEEL#2    timeout error on motion [src/filters/filters.c:2084]
375   2012/04/16   01:46:58.490407   FILTERS    WHEEL#1    timeout error on motion [src/filters/filters.c:2084] 
  • 02:00 UT Run testfiltersBR_dome again...
  • 02:20 UT blue filters hardware failure again... blue HK was off, graph shows intermittent readings.
  • 02:24 UT off/on

Start Transformation for AGw3

  • center of rotation: at Stone G field PA=0
  • luci.20120416.0004.fits (5-7 are 10 sec and 3 ndits)
  • PA=0, 8 is a 10sec by 3ndit int mode exposure, but this was in ACTIVE mode
  • PA=0, 9 is a 10 sec by 3 ndit int mode in track mode
  • Steve offset rotator by 90 deg so now PA=90, track mode
  • PA=90, image 10
  • Steve sends rotator to make PA=180, track mode, image 11
  • now to PA=270, track mode, image 12
  • now to PA=0, track mode, 13
  • back to PA=180, track mode, 14
  • dit=20, ndit=4, slew to hold rotator while integrating...preset cancelled, but still tracking at position and rotator
    • set rotator to 0, slew to hold 360... ellipse.
  • flexure compensation not on...
  • flexure compensation on now.
    • slew to hold rotator to 0 -- this is number 16.
  • but flex comp will not work when rotator is slewing...
  • send track preset PA=0, turn on flex comp on: 17, turn off flex comp. Lost collimation...
  • active preset to collimate at PA=0
  • track preset to PA=0 (alt 56), flex on, 18
  • track preset to PA=180, flex on, 19
  • track preset to PA=90, flex on, 20
  • track preset to PA=270, flex on, 21
  • track preset to PA=0, flex on, wait..., 22

  • Doug testing script with no instrument setup except for flex comp ON. Failed (due to corrupted database problem).
  • Edited script to remove this and add a pause...

  • 22:00 Starting on Stone_G, but setting... 52 to 40 deg.
    • flex comp on: 24, 25, 26, 27, 28, 29, 30 all at PA=0...
    • turn off flex compensation while sending preset which moves rotator
    • flex comp on. 31,...,34
    • flex comp on 35, 36, ...37
    • flex comp on ... 50 seeing is going off - around 2" now, at elev 46 degrees.
  • 23:15 Done collecting this set of data --- Doug offline to work on data for a couple of hours.

  • LBC: Blue HK was off. Turn off/on systems & HK and try testfiltersBR_onsky script. Red data taking didn't happen. Restarted script and getting red/blue data now.

  • 23:30 As we had finished the transf data collection at elev ~ 45 deg, and seeing had gotten so much worse from the start of data collection (finished at 2" from 0.8"), we slewed to higher (66 deg) elevation to check seeing. In fact it was bad there too, 1.8" and DIMM is 2.4". So it looks like the trend towards worse seeing was over time, rather than over elevation. Steve said wind direction had changed. * 23:40 At 78 deg elevation, first star was 2" but this one is at a similar elevation (74 deg) is 1.2". DIMM measurements correlate with GCS ones, so seeing appears really to be variable.

  • While computing transf matrix & rotator center, running LBC. 01:30 had a blue filter hardware failure.
  • Rotator center:
    • geomap on 10 stars matched between 0/180 and 90/270.
    • 0/180 frames 18/19 ---> 1024,1029
      • imexam to measure centroids of stars on each image.
      • make a table of:
x1_pa0 y1_pa0 x1_pa180 x2_pa180
x2_pa0 y2_pa0 x1_pa180 x2_pa180
... ... ... ...
x3_pa0 y3_pa0 x3_pa180 x3_pa180
      • input this table to geomap
      • record xshift, yshift that geomap outputs.
      • xshift/2 = rotcen_x and yshift/2 = rotcen_y.
    • 90/270 frames 20/21 ---> 1022, 1027.
    • (1023,1028) is rotator center I measure.

  • Closed for high humidity at 09:30 UT. We were not taking sky-data anyway --- awaiting transformation data, checking rotator center and testing LBC instead.
    • Used geomap on 6 matching stars on the images from 20120304 to check my procedure for measuring rotator center (based on mimicing the steps of LBC.rotcen by hand).
      • from luci.20120304.0395 (PA=0) vs luci.20120304.0394 (PA=180) I find (1010, 1049). This is close to what Dave reported in the wiki for these data http://wiki.lbto.org/twiki/bin/view/Instrumentation/RecalibrateLUCIAfterInstall. So it seems that the rotator center really has shifted to (1023, 1028).
      • I later did a similar check with the luci.20120304.0396 (PA=90) and luci.20120304.0397.fits (PA=270) and got (1013,1050), close again to what Doug got.

  • 11:15 Doug said that Jesper sent a transformation matrix.
    • He implemented it, and Steve restarted the AGW and GCS and OACserver
  • 11:30 Steve sending us to StoneK (low elev @ 41 deg but the only one available now).
    • starting to collect data at 11:42... first image will be 61
    • new pa... 65,...70
    • new pa 71...during 75 the guiding was not so good, because the star was too bright ---
    • new pa - didn't turn off flex comp during rotator move. but guider on 76, is good.
    • new pa...but no star found. Failed preset because too light. At 33 deg elevation.

  • 12:18 Closing.

-- OlgaKuhn - 16 Apr 2012
Topic revision: r6 - 16 Apr 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback